OpenStudy (anonymous):

need help solving the integral of x/(1+x^4)dx.

OpenStudy (anonymous):

\[\int\limits_{}^{}x/(1+x^4)dx\]

OpenStudy (amistre64):

x = tan^2 ?

OpenStudy (anonymous):

replace \[u=x^2\] get inverse tangent in one step

OpenStudy (amistre64):

thats the thought ... :)

OpenStudy (amistre64):

but then x = sqrt(u)

OpenStudy (anonymous):

that is if you remember that \[\frac{d}{dx}\tan^{-1}(x)=\frac{1}{1+x^2}\]

OpenStudy (dumbcow):

first use u = x^2 du = 2x dx then use trig substitution u = tan(theta) du = sec^2

OpenStudy (anonymous):

\[u=x^2\] \[du=2xdx\] \[\frac{1}{2}\int\frac{1}{1+u^2}du=\frac{1}{2}\tan^{-1}(u)\]]

OpenStudy (amistre64):

i sometimes forget to convert the dx:du part :)

OpenStudy (anonymous):

you can use a trig sub if you like, but if you already know that the derivative of arctangent is \[\frac{1}{1+x^2}\] then it is not necessary

OpenStudy (anonymous):

oh so "final answer" is \[\frac{1}{2}\tan^{-1}(x^2)\] and don't forget the stupid "plus cee"

OpenStudy (anonymous):

Thanks, everyone, for the help :) I had started it out, but it seemed too easy, so thought I'd double check. It's not a complicated prob at all...